# Generative Representational Learning of Foundation Models for Recommendation

### v1

@@ 1 Introduction
It has been a long-standing goal to develop a single foundation model to excel across diverse tasks in the field of artificial intelligence bommasani2021opportunities . Foundation models, characterized by their extensive knowledge bases and large-scale parameters, have demonstrated remarkable generalization capabilities across various domains, including natural language processing and computer vision. These models rely on large-scale pretraining to capture rich representations, enabling them to adapt to a wide range of downstream tasks with minimal finetuning muennighoff2025grit . Recently, the influence of foundation models has extended to the field of recommendation systems, where they offer unique opportunities to enhance personalization and prediction accuracy by effectively modeling complex user-item interactions and contextual information huang2024foundation of distinct sources and formats.
@@
First, existing approaches overlook the crucial embedding [^1] In this work, the term “embedding” is interchangeable with “representational”. tasks, which are fundamental to recommendation performance. Embedding techniques transform high-dimensional, discrete features such as user and item identifiers or descriptions into low-dimensional, continuous vectors, capturing intricate relationships and facilitating efficient computation. High-quality embeddings are essential for various components of recommendation systems, including retrieval, recall, and visualization liu2023embedding ; chen2022deep . Second, current models struggle with the complexities of multi-task learning, particularly in managing the knowledge sharing & conflict and the inconsistent convergence speeds among tasks. These issues can hinder the effective training of foundation models intended to perform well across diverse recommendation tasks, limiting their practical applicability in real-world systems wu2023robust .
@@ 2 Related Works
Early research on recommendation focused on optimizing single-task generative models hidasi2015gru4rec ; dong2017seq2seq ; wang2017irgan ; liang2018variational , such as producing personalized reviews via neural review generators dong2017neural_review_gen or generating item descriptions from user behavior sequences yang2018behavior_driven_desc . These models are typically trained from scratch and lack generalization across tasks. In the wake of large language models, recent works introduce foundation models tailored to e-commerce scenarios shen2023large ; zhou2024large ; wang2024unirec ; liu2023fim . LLaMA-E shi2024llamae tunes LLaMA grattafiori2024llama3herdmodels to support various product authoring tasks. EcomGPT li2024EcomGPT leverages generative pretraining over large-scale e-commerce corpora to unify tasks such as product Q&A, summarization, and marketing text generation. eCeLLM peng2024eCeLLM further develops a general-purpose e-commerce model by training on high-quality instruction datasets, enabling both in-domain and cross-task generalization. Such efforts demonstrate the growing interest in generative foundation models for recommendation. However, the field still lacks foundation models that support both representation and generative capabilities. To the best of our knowledge, we are the first to develop a unified generative representational learning framework for recommendation foundation models.
@@ 5.1 Experimental Setup
Our backbone LLM is Mistral-7B-Instruct jiang2023mistral7b . We finetune the backbone LLM on 8 $\times$ H100 GPUs for 3 epochs on the RecFound dataset, using a total batch size of 2048 for embedding tasks and 1024 for generative tasks. We use AdamW loshchilov2017decoupled with a learning rate of $2\!\times\!10^{-5}$. For TMoLE, we apply LoRA hu2021lora to the query/key/value/output projections with rank $r=16$, $\alpha=64$, and dropout rate 0.1. We instantiate $N=6$ low-rank experts for each projection, with the number of embedding experts $E=2$, generative experts $G=2$, and shared experts $S=2$. The task embedding size is 512. For S2Sched, we set the warmup ratio $\varsigma$ as 10%, the length of the history window $L$ as 64 (last steps), and temperature $\tau$ as 10. We uniformly sample 128 validation data instances to calculate the normalized validation loss for each task at each training step.

### v3

@@ 1 Introduction
It has been a long-standing goal to develop a single foundation model to excel across diverse tasks in the field of artificial intelligence bommasani2021opportunities . Foundation models, characterized by their extensive knowledge bases and large-scale parameters, have demonstrated remarkable generalization capabilities across various domains, including natural language processing and computer vision. These models rely on large-scale pretraining to capture rich representations, enabling them to adapt to a wide range of downstream tasks with minimal finetuning muennighoff2025grit . Recently, the influence of foundation models has extended to the field of recommendation systems, where they offer unique opportunities to enhance personalization and prediction accuracy by effectively modeling complex user-item interactions and contextual information of distinct sources and formats huang2024foundation ; lin2024rella ; lin2024clickprompt ; xi2024memocrs ; wang2024flip .
@@
First, existing approaches overlook the crucial embedding [^1] In this work, the term “embedding” is interchangeable with “representational”. tasks, which are fundamental to recommendation performance. Embedding techniques transform high-dimensional, discrete features such as user and item identifiers or descriptions into low-dimensional, continuous vectors, capturing intricate relationships and facilitating efficient computation. High-quality embeddings are essential for various components of recommendation systems, including retrieval, recall, and visualization huang2024comprehensive ; liu2023embedding ; chen2022deep . Second, current models struggle with the complexities of multi-task learning, particularly in managing the knowledge sharing & conflict and the inconsistent convergence speeds among tasks. These issues can hinder the effective training of foundation models intended to perform well across diverse recommendation tasks, limiting their practical applicability in real-world systems xi2024towards ; wu2023robust .
@@ 2 Related Works
Early research on recommendation focused on optimizing single-task generative models hidasi2015gru4rec ; dong2017seq2seq ; wang2017irgan ; liang2018variational ; lin2025can ; xi2024decoding , such as producing personalized reviews via neural review generators dong2017neural_review_gen or generating item descriptions from user behavior sequences xi2024efficient ; yang2018behavior_driven_desc . These models are typically trained from scratch and lack generalization across tasks. In the wake of large language models, recent works introduce foundation models tailored to e-commerce scenarios shen2023large ; zhou2024large ; wang2024unirec ; liu2023fim . LLaMA-E shi2024llamae tunes LLaMA grattafiori2024llama3herdmodels to support various product authoring tasks. EcomGPT li2024EcomGPT leverages generative pretraining over large-scale e-commerce corpora to unify tasks such as product Q&A, summarization, and marketing text generation. eCeLLM peng2024eCeLLM further develops a general-purpose e-commerce model by training on high-quality instruction datasets, enabling both in-domain and cross-task generalization. Such efforts demonstrate the growing interest in generative foundation models for recommendation. However, the field still lacks foundation models that support both representation and generative capabilities. To the best of our knowledge, we are the first to develop a unified generative representational learning framework for recommendation foundation models.
@@ 5.1 Experimental Setup
Our backbone LLM is Mistral-7B-Instruct jiang2023mistral7b . We finetune the backbone LLM on the RecFound dataset, using a total batch size of 2048 for embedding tasks and 1024 for generative tasks. We use AdamW loshchilov2017decoupled with a learning rate of $2\!\times\!10^{-5}$. For TMoLE, we apply LoRA hu2021lora to the query/key/value/output projections with rank $r=16$, $\alpha=64$, and dropout rate 0.1. We instantiate $N=6$ low-rank experts for each projection, with the number of embedding experts $E=2$, generative experts $G=2$, and shared experts $S=2$. The task embedding size is 512. For S2Sched, we set the warmup ratio $\varsigma$ as 10%, the length of the history window $L$ as 64 (last steps), and temperature $\tau$ as 10. We uniformly sample 128 validation data instances to calculate the normalized validation loss for each task at each training step.